OpenStudy (anonymous):

The town of Beeville is made up of two neighborhoods: Forest Ridge and Lakeland. Forest Ridge has a population of 11,098 and an area of 13 square miles. Lakeland has a population of 15,810 and an area of 18 square miles. What is the population density of Beeville? A 868 B 878 C 866 D 854

OpenStudy (anonymous):

@d3Xter

OpenStudy (anonymous):

okay take the pop of the town divide them by there miles square and add them together and divide by 2

OpenStudy (anonymous):

(11098/13 + 15810 / 18)/2=_____

OpenStudy (anonymous):

ya I got 486 something I dont think your right

OpenStudy (anonymous):

ok 11098 / 13 = then 15810/18= add then then divide by 2

OpenStudy (anonymous):

and them than / by 2

OpenStudy (anonymous):

still getting the same answer

OpenStudy (anonymous):

11098 / 13 =853 15810/18=878 853+878=1731 1731/2=

OpenStudy (phi):

I would do (total population) / (total area)

OpenStudy (anonymous):

865.5 I think I see what I was doing wrong now

OpenStudy (anonymous):

and if you round it wat do you get

OpenStudy (phi):

(11098 + 15810)/(13+18)